OpenStudy (anonymous):

f(x)=(x^2+10)(1-x^2) find interval of increasing, decreasing, concave up, concave down, local max and min and infection point

OpenStudy (anonymous):

Did you take the derivative?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What did you end up with?

OpenStudy (anonymous):

2(1-x^2)-2(x^2+10)

OpenStudy (anonymous):

That's not quite right by the product rule \[\frac{d}{dx}[(x^2+10)(1-x^2)] \] \[=(x^2+10)\frac{d}{dx} [(1-x^2)] + (1-x^2) \frac{d}{dx} [(x^2+10)]\] \[=-2x(x^2+10) + 2x(1-x^2) = 2x(1-x^2 -x^2 -10)\] \[=2x(-2x^2 -9) = -2x(2x^2 + 9)\]